1.     A driver of an ambulance going 70 km/h suddenly sees another car 35.0 m ahead. It takes the driver 0.62 s before he applies the brakes. Once he does begin to brake, he reduces speed at a rate of 13.0 m/s².

 

1.1  Determine if the driver will hit the car or not.                                               (6)

 

What would be the maximum speed at which the ambulance could travel and NOT hit the car 35.0m ahead?             


1
Expert's answer
2021-06-01T10:51:39-0400

"a=\frac{v-v_0}{t}=13" m/s2

Time of breaking:

"t=\frac{v-v_0}{a}=70/(3.6\cdot13)=1.5\ s"

Breaking distance:

"d=70\cdot0.62/3.6+v_0t-at^2/2"

"d=70\cdot0.62/3.6+70\cdot1.5/3.6-13\cdot1.5^2/2=26.6\ m"

Since distance d < 35 m, the driver will not hit the car.


For maximum speed:

"d=35=v_{max}\cdot0.62+v_{max}t-at^2/2"

"d=35=v_{max}\cdot0.62+v_{max}^2/a-v^2_{max}/(2a)=v_{max}\cdot0.62+v^2_{max}/(2a)"

"v^2_{max}+16.12v_{max}-910=0"


"v_{max}=\frac{-16.12+\sqrt{16.12^2+4\cdot910}}{2}=23.16\ m/s=83.4\ km/h"